OpenStudy (anonymous):

Give the values of a, b, and c needed to write the equation's general form. 2/3(x-4)(x+5)=1 A = 2/3; B = 1; C = -20 A = 2; B = 2; C = 43 A = 2; B = 2; C = -43

OpenStudy (mertsj):

\[\frac{2}{3}(x^2+x-20)=1\] \[2x^2+2x-40=3\] \[2x^2+2x-43=0\] a=2, b=2, c=-43